OpenStudy (anonymous):

Create your own binomial expression with a radical in the second term. Part 1: Identify its conjugate and explain, in complete sentences, why it is the conjugate. (1 point) Part 2: Multiply your original binomial expression and its conjugate. What happened to the radicals and why?

OpenStudy (anonymous):

how about \(2+\sqrt5\) or \(3+\sqrt7\) or \(4+\sqrt{10}\) ?

OpenStudy (anonymous):

Yeah that's good but why is it conjugate ? I'm having trouble with this chapter ugh -__- difficult trying to understand

OpenStudy (anonymous):

the conjugate of \(a+\sqrt{b}\) is \(a-\sqrt{b}\)

OpenStudy (anonymous):

so for example the conjugates of each one i wrote above, replace the \(+\) by a \(-\)

OpenStudy (anonymous):

Ohh okay i see now.

OpenStudy (anonymous):

so it becomes \[2-\sqrt{5}\] ?